A 12 metre pole casts a shadow 10 metres long. If measurements are taken at the same time:
i) How tall is a tree, if its shadow is 25.5 metres long?
ii) How long is a shadow, if the building is 110 metres high?

Answer by Alan3354(69443)

A 12 metre pole casts a shadow 10 metres long. If measurements are taken at the same time:
i) How tall is a tree, if its shadow is 25.5 metres long?
The ratio is the same.
height/shadow = 12/10
---------
12/10 = h/25.5
10h = 306
h = 30.6 m
=================
ii) How long is a shadow, if the building is 110 metres high?
12/10 = 110/s